OpenStudy (anonymous):

Find a ⋅ b. a = 9i + 5j, b = 4i + 3j

OpenStudy (anonymous):

51 <13, 8> <36, 15> -21

OpenStudy (anonymous):

@sauravshakya @A.Avinash_Goutham @amistre64

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Its a dot product

OpenStudy (anonymous):

Hint: if a=xi+yj AND b=x'i+y'j THEN a . b = x*x' + y*y'

OpenStudy (anonymous):

can you show me how to work out this problem.

OpenStudy (anonymous):

First can u tell me the values of x , x' ,y , y'

OpenStudy (anonymous):

Just compare what I gave in the Hint and your question